# Unexpected consequences of Post-Quantum theories in the graph-theoretical approach to correlations

### v1

@@ II Results
By Proposition 1, the largest theory allowed by EP to the experiment is $Y(G)=\mathrm{abl\>}\mathrm{Q(\bar{G})}$. As shown in Ref. [33], for any experiment the equality $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$ holds. Consequently, $\mathrm{abl\>Q(\bar{G})}=\mathrm{abl\>abl\>Q(G)}$, and since both $\mathrm{Q(G)}$ and $\mathrm{Q(\bar{G})}$ are convex corners, it follows that $Y(G)=\mathrm{abl\>Q(\bar{G})}=\mathrm{Q(G)}$. ∎
Moreover, we can see Proposition 1 as a generalization of Result $1$ of Ref. [22] framed in a theory-independent manner, as it determines the largest set allowed by EP for the experiment, $Y(G)=\mathrm{abl}\>X(\bar{G})$, for any assumed theory describing the complementary experiment, $X(\bar{G})$, ranging from the noncontextual to the $E_{1}$ theory.
An intriguing fact arises from the result $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$ used in the previous proof. In essence, the theory assumed for the complementary experiment, quantum, is precisely the same theory singled out by EP. This stands in contrast to the case where classical or $E_{1}$ theory were assumed (recall that if classical theory were the sole framework for describing physical phenomena in a complementary experiment, EP, by singling out $E_{1}$ theory for the experiment, asserts that the latter is governed by a completely different set of rules— or, in other words, distinct physical laws). That a theory is consistently utilized to describe each and every experiment is certainly a desirable property, for easy-to-understand reasons. In this context, the following question arises: are there other theories $U$ that satisfy the self-duality equation $\mathrm{U(\bar{G})}=\mathrm{abl\>U(G)}$? Should this be the case, Proposition 1 would yield the same conclusions as those drawn from the quantum scenario. Furthermore, as far as EP can ascertain, both theories would be equally valid in describing empirical data. We claim that if the uniqueness of QT as the solution to the self-duality equation is demonstrated, then it is shown EP explains QT within the CSW framework. A partial answer to the above question is now discussed.
It seems adequate to argue that for such hypothetical theory, $U$, to be of physical interest and potentially lead to new physics—for instance, by allowing the possibility of violating quantum bounds of NC inequalities—it must extend beyond QT while still encompassing all of its predictions. In the following, we argue that if such a theory exists, it would imply that certain quantum behaviors are prohibited in an independent experiment.
@@
If every quantum behavior is experimentally accessible, then EP excludes post-quantum behaviors. Equivalently, quantum theory and EP exclude post-quantum theories.
@@
We will prove the contrapositive. That is, if the set of behaviors (or the theory) associated with the experiment $\bar{G}$ is post-quantum, then this implies that the EP restricts the set of behaviors (or the theory) for the experiment $G$, such that it becomes strictly contained within the set of quantum behaviors. In other words, under these conditions, certain quantum behaviors would become experimentally inaccessible. Let then be $U(\bar{G})\equiv\mathrm{Q(\bar{G})}\cup W(\bar{G})$ the set behaviors for the experiment $\bar{G}$, where each $\omega^{\prime}\in W(\bar{G})$ corresponds to a post-quantum behaviour; i.e., $\omega^{\prime}\in\mathrm{E_{1}(\bar{G})}\setminus\mathrm{Q(\bar{G})}$. We assume that $U(\bar{G})$ is a convex corner that satisfies the self-duality equation $U(\bar{G})=\mathrm{abl\>}U(G)$ (though the main result still holds even in the absence of latter assumption; no further assumptions are made regarding $W(\bar{G})$, other than its non-emptiness). Thus, our post-quantum theory is represented by $U(\bar{G})$. From Proposition 1, the theory singled out by EP to describe the experiment is given by $Y(G)=\mathrm{abl\>}U(\bar{G})=\mathrm{abl\>}\{Q(\bar{G})\cup W(\bar{G})\}$. More explicitly,
$Y(G)=\Big{\{}p\geq 0|\sum_{i}p_{i}p_{i}^{\prime}\leq 1\\ \>\forall\>p^{\prime}\in U(\bar{G})=\mathrm{Q(\bar{G})\cup W(\bar{G})}\Big{\}}.$ (6)
Therefore,
$Y(G)=\Big{\{}p\geq 0\>|\>\big{[}\ \sum_{i}p_{i}q_{i}^{\prime}\leq 1\>\forall\>q^{\prime}\in\mathrm{Q(\bar{G})}\big{]}\ \\ \land\big{[}\ \sum_{i}p_{i}w_{i}^{\prime}\leq 1\>\forall\>w^{\prime}\in W(\bar{G})\big{]}\ \Big{\}},$ (7)
and finally,
$Y(G)=\mathrm{Q(G)}\cap\Big{\{}p\geq 0\>|\>\sum_{i}p_{i}w_{i}^{\prime}\leq 1\>\forall\>w^{\prime}\in W(\bar{G})\Big{\}}.$ (8)
From Eq. (8), it follows immediately that $Y(G)\subseteq\mathrm{Q(G)}$. Using the self-duality property of QT, we have
$\mathrm{Q(\bar{G})}=\Big{\{}p^{\prime}(\bar{G})\geq 0|\sum_{i}p_{i}^{\prime}p_{i}\leq 1\>\forall\>p\in\mathrm{Q(G)}\Big{\}};$ (9)
for any $p^{\prime}(\bar{G})\in\mathrm{E_{1}(\bar{G})\setminus Q(\bar{G})}$, there exists $p\in\mathrm{Q(G)}$ such that
$\sum_{i}p^{\prime}_{i}p_{i}>1.$ (10)
Let $p^{\prime}(\bar{G})=w^{\prime}\in W(\bar{G})$; then there exists $p\in\mathrm{Q(G)}$ such that
@@
Any such $p(G)$ is excluded from the definition of $Y(G)$ in Eq. (8) and, as a matter of fact, $Y(G)=\mathrm{abl\>U(\bar{G})}\subset\mathrm{Q(G)}$. It is even possible to show that all excluded behaviors are post-classical (or, if you will, genuinely quantum), that is, $p(G)\in\mathrm{Q(G)\setminus NC(G)}$. Indeed, from the duality between classical and $E_{1}$ theory used in corollary 1, we have
@@
As a consequence, for all $w^{\prime}\in W(\bar{G})$, any $p(G)\in\mathrm{NC(G)}$ satisfies $\sum_{i}p_{i}w_{i}\leq 1$.
Any Post-Quantum theory assumed for the complementary experiment leads to a theory strictly smaller than Quantum for the experiment, singled out by EP. ∎
Let’s now explore the implications of our main result. Under the assumptions that EP holds and Post-Quantum theories are allowed for at least one complementary experiment, our findings ensure that some genuinely quantum behaviors become prohibited. However, quantum theory has been extensively tested in various experimental contexts, and its predictions have been found to agree with observation across an extraordinarily wide range of scenarios [40, 41, 42, 43, 1, 2, 3]. While certain quantum behaviors may be technically challenging to realize [44, 45, 46, 47], there is no fundamental reason to believe that any of them are impossible to achieve in Nature. Therefore, it is plausible to argue that one of the two assumptions must be false.
@@
Given that EP is our chosen physical principle, we will first assume that EP is indeed respected in Nature—a view supported by experimental evidence [19]. Consequently, we conclude that EP imposes a fundamental limitation, ruling out the existence of theories that extend beyond Quantum Theory. This conclusion generalizes ATC’s second result in [22] to arbitrary experiments, beyond the specific case of self-complementary graphs.
@@
Continuing with the assumption that no quantum behavior is fundamentally impossible to realize in Nature, an alternative perspective for those who do not view QT as the ultimate theory of Nature is that the EP does not hold universally. In other words, if a Post-Quantum behavior were ever observed in an experiment, our result would ensure a breakdown of the Exclusivity Principle.
Finally, the question of whether there exist theories $U(\bar{G})$ that satisfy the self-duality equation can now be interpreted as an enhancement of our main result. Indeed, if it can be proven that QT is the only theory with such a property, then any alternative theory would fail to be universal, as some of its predictions would be prohibited in independent experiments. This would demonstrate that EP uniquely singles out QT, without presupposing that all quantum behaviors are necessarily realizable in Nature. Moreover, our result advances the resolution of this problem by showing that a theory satisfying the self-duality equation can neither strictly contain quantum theory nor be strictly contained by it.
@@ III Conclusion
In this work, we extended previous results on the quantum correlations problem within the CSW framework. Specifically, we examined a type of well-known composite correlation experiment, Yan’s construction. We identified that the largest theory allowed by the Exclusivity Principle (EP) to describe an experiment, given a theory for the complementary experiment, aligns with the mathematical structure of anti-blocking duality. This insight enabled us to determine, for any assumed theory of any complementary experiment, the corresponding theory selected by EP for the original experiment (Proposition 1). We then explored the implications of this result for classical and $E_{1}$ theories (Corollary 1), and more importantly, for Quantum Theory (Remark 2), where we recovered a previously established result from the literature.
Building on Proposition 1 we derived our main result, where we demonstrated that Quantum Theory and EP forbid Post-Quantum theories to produce results distinct from the quantum ones. In other words, by assuming that EP holds, we derive a fundamental limitation: EP excludes the possibility of theories extending beyond Quantum Theory, generalizing previous results to arbitrary experiments. Alternatively, if Post-Quantum behaviors are observed and the corresponding Sub-Quantum counterpart cannot be certified, this would indicate that EP is not a valid universal principle.
Finally, we propose that another way to show that EP uniquely selects QT for any experiment is by proving that QT is the unique solution to the self-duality equation $U(\bar{G})=\mathrm{abl\>U(G)}$. Our work also serves as a bridge between the findings of ATC’s 2014 work [22] (along with earlier works that follow the same logical structure), by generalizing some of its results, and Cabello’s 2019 work [26], by reaching a similar conclusion for the quantum correlations problem. While both approaches highlight the role of EP, our approach reaches this result through Yan’s more streamlined construction, offering a simpler framework.
@@ Acknowledgements.
We thank Bárbara Amaral, Adán Cabello and Rafael Rabelo for enlightening discussions. This study was financed in part by the Brazilian National Council for Scientific and Technological Development (CNPq) and in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. This work was also supported by CNPq under grant no. 311314/2023-6, and it is part of the Brazilian National Institute for Science and Technology in Quantum Information (INCT-QI).

### v2

@@ II Results
By Proposition 1, the largest correlation set allowed by EP to the experiment $\mathrm{G}$ is $\mathrm{Y(G)}=\mathrm{abl\>}\mathrm{Q(\bar{G})}$. As shown in Ref. [Knuth1994], for any experiment $\mathrm{G}$ the equality $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$ holds. Consequently, $\mathrm{abl\>Q(\bar{G})}=\mathrm{abl\>abl\>Q(G)}=\mathrm{Q(G)}$, since both $\mathrm{Q(G)}$ and $\mathrm{Q(\bar{G})}$ are convex corners. Thus, it follows that $\mathrm{Y(G)}=\mathrm{abl\>Q(\bar{G})}=\mathrm{Q(G)}$. ∎
Moreover, we can see Proposition 1 as a generalization of Result $1$ of Ref. [Amaral2014] framed in a theory-independent manner, as it determines the largest set allowed by the EP for the experiment, $\mathrm{Y(G)}=\mathrm{abl}\>\mathrm{X(\bar{G})}$, for any assumed correlation set describing the complementary experiment, $\mathrm{X(\bar{G})}$, ranging from the classical to the $E_{1}$ correlations sets.
We now present our main result, which follows from allowing the complementary experiment to be described by some post-quantum behavior. Interestingly, in this case, the EP leads to an unexpected and somewhat drastic consequence for the description of the experiment itself.
@@
Assuming that the EP holds and that post-quantum behaviors are accessible in at least one experiment, it follows that certain genuinely quantum behaviors become forbidden in a related yet completely independent experiment. More precisely, if a post-quantum behaviour is accessible in an experiment $\mathrm{\bar{G}}$, then EP implies that some quantum behavior will be missed in the experiment $\mathrm{G}$.
@@
Let $U(\bar{G})$ be the set of behaviors for the experiment $\bar{G}$, where there exist $w^{\prime}\in U(\bar{G})$ corresponding to a post-quantum behaviour; i.e., $w^{\prime}\in\mathrm{E_{1}(\bar{G})}\setminus\mathrm{Q(\bar{G})}$. From Proposition 1, the set of correlations singled out by the EP to describe the experiment is given by $\mathrm{Y(G)}=\mathrm{abl\>}U(\bar{G})$. More explicitly,
$\mathrm{Y(G)}=\Big\{p\geq 0|\sum_{i}p_{i}p_{i}^{\prime}\leq 1\>\forall\>p^{\prime}\in U(\bar{G})\Big\}.$ (6)
Using the self-duality property of QT, we have
$\mathrm{Q(\bar{G})}=\Big\{p^{\prime}\geq 0|\sum_{i}p_{i}^{\prime}p_{i}\leq 1\>\forall\>p\in\mathrm{Q(G)}\Big\};$ (7)
thus, for any $\tilde{p}\in\mathrm{E_{1}(\bar{G})\setminus Q(\bar{G})}$, there exists $p\in\mathrm{Q(G)}$ such that
$\sum_{i}\tilde{p}_{i}p_{i}>1.$ (8)
Let $\tilde{p}=w^{\prime}$; then there exists $p\in\mathrm{Q(G)}$ such that
@@
Any such $p$ is excluded from the definition of $\mathrm{Y(G)}$ in Eq. (6) and, as a matter of fact, $\mathrm{Y(G)}=\mathrm{abl\>U(\bar{G})}\nsupseteq\mathrm{Q(G)}$. It is even possible to show that all excluded behaviors are post-classical (or, if you will, genuinely quantum), that is, $p\in\mathrm{Q(G)\setminus NC(G)}$. Indeed, from the duality between classical and $E_{1}$ theory used in corollary 1, we have
@@
As a consequence, for all post-quantum $w^{\prime}\in U(\bar{G})$, any $p\in\mathrm{NC(G)}$ satisfies $\sum_{i}p_{i}w_{i}\leq 1$. ∎
Let’s now explore the implications of our main result. Under the assumptions that the EP holds and some post-quantum correlation is allowed for at least one complementary experiment, our findings ensure that some genuinely quantum behaviors become prohibited. However, quantum theory has been extensively tested in various experimental contexts, and its predictions have been found to agree with observation across an extraordinarily wide range of scenarios [bouwmeester_experimental_1997, razavy_quantum_2003, hanneke_new_2008, juffmann_real-time_2012, Giustina2015, Hensen2015, Shalm2015]. While certain quantum behaviors may be technically challenging to realize [nadlinger_experimental_2022, zhao_loophole-free_2024, bernstein_quantum_1997, watrous2008quantumcomputationalcomplexity], there is no fundamental reason to believe that any of them is impossible to achieve in Nature. In this way, it seems adequate to assume that every quantum behavior is, in principle, realizable in Nature. Granted, this assumption cannot be empirically verified, as it would require certifying an infinite number of quantum correlations. Nevertheless, it has been explicitly adopted in some works [dirac_principles_2010], based on the premise that every quantum state and measurement predicted by quantum theory can be realized in Nature—a similar assumption, known as the no-restriction hypothesis, is also commonly adopted in the context of GPTs [Chiribella2010, Chiribella2011]. Moreover, we argue that an analogous assumption is implicitly made in many other works [braunstein_wringing_1990, BBDH1997PRL, Cerf_GHZParadoxes_2002, acin_randomness_2012, araujo_all_2013, ZhaoQ2023, vieira2024test]. For instance, it underlies studies in which families of states and measurements—sometimes highly intricate—are assumed, theoretical results are derived from them, and these results are then regarded as valid despite the lack of direct empirical verification. To summarise, since our main result contradicts a premise we firmly uphold (that every quantum behavior is, in principle, realizable in Nature), it is plausible to argue that one of the other two assumptions (that the EP holds and that some experiment might be described by a post-quantum behavior) must be false.
@@
Given that the EP is our chosen physical principle, we will first assume that it is indeed respected in Nature—a view supported by experimental evidence [nawareg2013bounding]. Consequently, we conclude that the EP imposes a fundamental limitation, ruling out the existence of correlation sets that extend beyond the quantum one. This conclusion generalizes ATC’s second result in [Amaral2014] to arbitrary experiments, beyond the specific case of self-complementary graphs. At this point, a remark is in order. Our main conclusion is that, given our assumptions, post-quantum correlations are ruled out. While this result primarily concerns correlation sets, it also has indirect implications for post-quantum theories. By a theory, we mean any set of rules that for any given graph $\mathrm{G}$ associates a set of behaviors $\mathrm{U(G)}$, for instance GPTs within the CSW framework. Such a theory will be post-quantum if there exists at least one graph $\mathrm{G}$, such that $\mathrm{U(G)}\nsubseteq\mathrm{Q(G)}$. Since our results exclude the realization of any such correlation set, it is possible to argue that they also exclude any theory capable of producing it.
@@
Continuing with the assumption that no quantum behavior is fundamentally impossible to realize in Nature, an alternative perspective for those who do not view QT as the ultimate theory of Nature is that the EP does not hold universally. In other words, if a Post-Quantum behavior were ever observed in an experiment, our result would ensure a breakdown of the Exclusivity Principle. However, note that discussing the EP in these post-quantum theories is meaningful only if a notion of sharpness for measurements is assumed within them. This notion of sharpness in Generalized Probabilistic Theories, along with the EP and their relationship, is discussed in [chiribella_measurement_2014, gonda_almost_2018].
Before concluding this section, we make a brief technical remark. In the proof of our main result, nothing was assumed about the exact form of the correlation set $U(\bar{G})$, other than the existence of some post-quantum behavior in it. Nonetheless, based on the previous discussion, it is only natural to argue that the most adequate form for us to assume for said correlation set is actually $U(\bar{G})\equiv\mathrm{Q(\bar{G})}\cup W(\bar{G})$, where each $w^{\prime}\in W(\bar{G})$ corresponds to a post-quantum behaviour; i.e., $w^{\prime}\in\mathrm{E_{1}(\bar{G})}\setminus\mathrm{Q(\bar{G})}$. That is, it extends beyond quantum correlations while still containing all of them. In fact, such a form of correlation set seems to be physically interesting, since it is consistent with current empirical data (we are able to obtain quantum bounds of NC inequalities), but might potentially lead to new physics by also allowing the possibility of violating said quantum bounds for some NC inequality. In this particular case, it is straightforward to see Proposition 1 yields
$\mathrm{Y(G)}=\mathrm{Q(G)}\cap\Big\{p\geq 0\>|\>\sum_{i}p_{i}w_{i}^{\prime}\leq 1\>\forall\>w^{\prime}\in W(\bar{G})\Big\},$ (11)
and it follows immediately that $\mathrm{Y(G)}\subseteq\mathrm{Q(G)}$ (that is, the EP constrains the correlation set of $G$ to be at most quantum). But, as shown before, the existence of post-quantum behaviors in the complementary experiment implies that some genuinely quantum behaviors of the experiment become forbidden, and thus $\mathrm{Y(G)}=\mathrm{abl\>U(\bar{G})}\subset\mathrm{Q(G)}$. In words, any post-quantum correlation set (containing the quantum) assumed for the complementary experiment leads to a correlation set strictly smaller than quantum for the experiment, singled out by the EP. To close this comment, note that even though some genuinely quantum behaviors of the experiment become forbidden, this is not the case for all of them; that is, $\mathrm{Y(G)}=\mathrm{abl\>U(\bar{G})}$, is, in general, a post-classical correlation set, containing $\mathrm{NC(G)}$ as a proper subset. In fact, recalling Corollary 1, $\mathrm{abl\>U(\bar{G})}$ is purely classical only if we choose $\mathrm{U(\bar{G})=Q(\bar{G})}\cup W(\bar{G})=\mathrm{E_{1}(\bar{G})}$.
@@ III Conclusion
In this work, we extended previous results on the quantum correlations problem within the CSW framework. Specifically, we examined a type of well-known composite correlation experiment, Yan’s construction. We identified that the largest correlation set allowed by the Exclusivity Principle (EP) to describe an experiment, given a correlation set for the complementary experiment, aligns with the mathematical structure of anti-blocking sets. This insight enabled us to determine, for any assumed correlation set of any complementary experiment, the corresponding correlation set selected by the EP for the original experiment (Proposition 1). We then explored the implications of this result for classical and $E_{1}$ sets (Corollary 1), and more importantly, for quantum correlations (Remark 2), where we recovered a previously established result from the literature.
Building on Proposition 1 we derived our main result, where we demonstrated that if Nature is at least quantum, then the EP forbids post-quantum behaviors in Nature. In other words, by assuming that every quantum behavior is, in principle, realizable in Nature and that the EP holds, we derive a fundamental limitation: the EP excludes the possibility of theories extending beyond Quantum Theory, generalizing previous results to arbitrary experiments. Alternatively, if Post-Quantum behaviors are observed, this would indicate that EP is not a valid universal principle. Our work also serves as a bridge between the findings of ATC’s 2014 work [Amaral2014] (along with earlier works that follow the same logical structure), by generalizing some of its results, and Cabello’s 2019 work [Cabello2019], by reaching a similar conclusion for the quantum correlations problem. While both approaches highlight the role of the EP, our approach reaches this result through Yan’s more streamlined construction, offering a simpler framework.
It would be interesting to compare our approach with other frameworks that aim to recover quantum theory from general principles. In particular, our results invite a comparison with axiomatic reconstructions developed within the GPT framework, especially in the context of the CSW approach [chiribella_measurement_2014, gonda_almost_2018]. Of special interest is the role of the No-Restriction Hypothesis (NRH) [Chiribella2010, Chiribella2011], which appears to bear some unveiled relation to the Exclusivity Principle.
To conclude this paper, let us mention a rather interesting question that aligns well with the general ideas presented here. An intriguing observation follows from the self-duality property $\mathrm{Q(\bar{G})}=\mathrm{abl\>Q(G)}$, which QT satisfies. In essence, the theory assumed for the complementary experiment—quantum—is precisely the same as the theory singled out by the EP for the original experiment (see Remark’s 2 proof). In this context, a natural question arises: are there other GPTs within the CSW framework whose associated correlation sets $U(.)$ satisfy the self-duality relation $\mathrm{U(\bar{G})}=\mathrm{abl\>U(G)}$? If so, Proposition 1 would yield conclusions analogous to those derived in the quantum case. If not, QT [^14] More precisely, the equivalence class of all theories that within the framework of CSW generate the same set of correlations as quantum theory for every graph G. is uniquely characterized by this self-duality property. Our main result contributes to addressing this issue by showing that any theory satisfying the self-duality relation can neither strictly contain QT nor be strictly contained by it.
@@ Acknowledgements.
We thank Bárbara Amaral, Adán Cabello and Rafael Rabelo for enlightening discussions. This study was financed in part by the Brazilian National Council for Scientific and Technological Development (CNPq) and in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. This work was also supported by CNPq under grants no. 311314/2023-6 and 445328/2024-0, and it is part of the Brazilian National Institute for Science and Technology in Quantum Information (INCT-QI).